OpenStudy (anonymous):

2x-3/x+4 >or equal to -5

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

need to solve for x

jhonyy9 (jhonyy9):

2x-3 ---- >= -5 x+4 2x-3>= -5(x+4) 2x-3>= -5x-20 2x+5x >= -20+3 7x >= -17 x >= -17/7